OpenStudy (anonymous):

PLS HELP MEH!!!!! The sum of two numbers is 37 . The smaller number is 19 less than the larger number. What are the numbers?

OpenStudy (anonymous):

\[x+y=37\\ x-y=19\]

OpenStudy (trojanpoem):

x + y =37 x > y y is 19 less than x so y = x- 19 x - 19 + x = 37 2x = 56 x = 28